# Exploring baby Julia sets in parameter space slices for Generalized McMullen Maps

### v1

@@ Abstract.
11footnotetext: 2010 MSC: Primary: 37F10, 32A20; Secondary: 32A19. Keywords: Complex Dynamical Systems, Mandelbrot, Polynomial-Like, Iteration, Douady, Hubbard
@@ 2. Preliminaries
Notation. We use $\mathcal{N}(S)$ to denote a small neighborhood (of unspecified size) of a set $S$.
@@ 2.1. Polynomial-Like Maps
Each polynomial-like map of degree $d$ has exactly d1d-1 critical points. Here we are only concerned with polynomial-like maps of degree two. A polynomial-like map of degree two is a two-to-one map from $U^{\prime}$ to $U$, and has a unique critical point in $U^{\prime}$. The “straightening theorem” shows this is an apt name.
@@
Now since $d=2$ in our case of interest, by the straightening theorem, we can associate each member $f_{\lambda}$ of an analytic family of polynomial-like maps of degree $2$ with a polynomial $P_{\chi(\lambda)}=z^{2}+\chi(\lambda)$ and a homeomorphism between neighborhoods of their Julia sets
@@
where $\mathcal{N}(K(f_{\lambda}|_{U^{\prime}_{\lambda}}))\subset U^{\prime}_{\lambda}$, and $\phi_{\lambda}$ is a hybrid equivalence from $f_{\lambda}$ to $P_{\chi(\lambda)}$ (conformal on filled Julia sets, quasi-conformal in neighborhoods), with
@@
We can improve upon this using McMullen and Sullivan’s [24]. In their Chp. 7, they describe the structural stability extended to the Riemann sphere, $\hat{\mathbb{C}}$. Recall a holomorphic family of rational maps $f_{\lambda}$ over a complex manifold $X$ is a holomorphic map $X\times\hat{\mathbb{C}}\to\hat{\mathbb{C}}$, given by $(\lambda,z)\mapsto f_{\lambda}(z).$ In a holomorphic family of rational maps, $f_{\lambda_{0}}$ is topologically stable if for all $\lambda$ in some neighborhood of $\lambda_{0}$, $f_{\lambda}$ and $f_{\lambda_{0}}$ are topologically conjugate (on the sphere). They show this implies there is in fact a quasi-conformal conjugacy (Theorem 7.1, and moreover the set of critical orbit relations is constant).
@@
They use the Harmonic $\lambda$-lemma of Bers-Royden (see [26] and [1]), which shows a holomorphic motion of a set $A$ in a neighborhood in parameter space can be extended uniquely to a holomorphic motion of the entire sphere, over a (smaller) neighborhood in parameter space (and the motion and its extension agree on the domain overlap). The uniqueness is what allows them to guarantee the compatibility with the dynamics.
In our setting, the compatibility with the dynamics implies that the $\tau$’s of Proposition 2.6 and Theorem 2.8 coincide, though Douady and Hubbard’s $\tau_{\lambda}$ was only defined near the Julia sets. (Of course, that’s why we used the same symbol for both!)
Moreover, McMullen-Sullivan apply their results to hyperbolic quadratic polynomials. Let $H\subset\mathbb{C}$ be the open set of hyperbolic parameters. Partition $H$ by setting $H(\infty)$ to the set of parameters whose critical orbits escape to $\infty$; $H_{0}(p)$ is the set of parameters whose critical point is periodic of order $p$; and $H(p)$ is the set of parameters whose critical orbit is attracted to a cycle of period $p$ (but strictly, $H_{0}(p)$ is not in $H(p)$).
[…]
@@
Finally, in this subsection we give Douady and Hubbard’s criteria for the existence of baby Mandelbrot sets in parameter space, as stated by Devaney based on Douady and Hubbard:
@@ 2.2. Preliminaries Regarding $R_{n,a,b}$
[4] For the family of functions $R_{n,a,b}$, let $\psi=\operatorname{Arg}(a)$ and for each $k=0,\ldots,2n-1,$ define $U_{a,k}^{\prime}$ (which also depends on $n$), as the polar rectangle:
@@
So $U^{\prime}_{a,k}$ is the polar rectangle with arguments centered at $(\psi+2k\pi)/2n$ and then plus or minus $\pi/2n$. Note $U^{\prime}_{a,k}$ does not depend on $b$, but of course its image does.
@@
Note before the shift and rotation, the ellipse $\mathcal{E}$ is centered at 0, its semi-major axis lies along the $x$-axis and has length $\left(2^{n}+{|a|}/{2}\right)$, and its semi-minor axis lies along the $y$-axis and has length $\left(2^{n}-{|a|}/{2}\right)$.
@@
[4] For each $k$ in $\{0,1,\ldots,2n-1\}$, if $k$ is even $U_{a}^{+}=R_{n,a,b}(U^{\prime}_{a,k})$ and this set is one half of the ellipse $\mathcal{E}$, including the minor axis and the critical value $v_{+}$. Moreover, $R_{n,a,b}$ maps $U^{\prime}_{a,k}$ $2:1$ onto $U_{a}^{+}$ for each even $k$.
@@ 3. Dynamical Plane Results for $r_{n,a}$
Excluding a “degenerate” case. For $n\geq 3$, if $a$ is small it can happen that $D_{-}=D_{+}$, that is, $v_{-}$ and $v_{+}$ are in the same (bounded) Fatou component. In that case, there is only one Fatou component, which the reader may imagine as a topological disk with an infinite tree of disks removed (the Fatou component containing $z=0$ and all of its preimages). Xiao, Qiu, and Yin ([27]) call such a $J$ a Cantor set of bubbles. In our subfamily with one critical point fixed, the hyperbolic component containing $a=0$ is a quite small teardrop-like shape aligned along the real axis, roughly in the bounds from $\Re(a)=-0.004$ to $\Re(a)=0.0025$ and $\Im(a)=-0.0025$ to $\Im(a)=0.0025.$ In the images in Figure 4 it is not really visible as it lies just to the right of the “deformed” baby $\mathcal{M}$-like set near the cusp of the cardioid shape of the bifurcation locus. But in the general, parameter space maps with $D_{-}=D_{+}$ are quite easy to find. For example, on the left of Figure 6 is the parameter slice which is the $b$-plane, holding $a$ small and constant, in this case for $n=4,a=0.0001+0.000i.$ The image appears like the parameter space for $z^{4}+b$, except with “bites” taken out of it due to the singular perturbation of 0 mapping to $\infty$. Any point in the main central annulus-like black component has both critical values in this common Fatou component. One such example is in this figure on the right, for $n=3,a=0.001+0.001i,c=0.25+0.4i$, although the situation for $n\geq 3$ is similar. The critical values are marked, and the shading inside of the filled Julia set illustrates the rate of attraction to the attracting fixed point.
@@ 4. Locating the bifurcation locus of $r_{n,a}$
We next show the bifurcation locus lies within an annulus. In the limit of the above cardioid as $n\to\infty$, the cardioid has cusp at $a(\pi)=0$, and max on the $x$-axis of $a(0)=1/4$, so is contained in an annulus centered at $1/8$, with a small inner radius <1/8<1/8, and outer radius about $1/7$. We show below is that the boundedness locus is contained in the annulus $\mathbb{A}(l,u)+1/8$, with center $1/8$, outer radius any $u>3/8$ and inner radius any $l\leq 1/32,$ for $n$ sufficiently large.
@@
One way to improve on Lemma 4.2 would be to show:
@@
The outline of a proof could be somewhat similar to the proof of Theorem 2 of [4] (or a simpler case in a lemma in [6]): show that inside the annulus guaranteed by Lemma 4.2, but outside of $\mathcal{N}_{\varepsilon}(\mathcal{S}_{n}),$ $|v_{-}|$ is bounded away from $1$, then apply Corollary 2.10. The fact that the spine depends on $n$ is an added complication that does not appear in [4], but computer images suggest $M_{n}(r_{n,a})\to\mathcal{S}_{\infty}$ as $n\to\infty$.
@@ 5. Centers of Principal Components in the $r_{n,a}$ Parameter Space
This gives us one trivial solution: If $a=0$, then $b-2\sqrt{a}=b+2\sqrt{a}=a^{1/2n}=0$. For the rest we need $v_{-}$ to be equal to a non-canonical choice of root.
@@
Around these 2n12n-1 points, we see two patterns emerge. Parameters defined by the above with $k$ even lie inside the disks we see in the bifurcation locus. The $a_{k}$ for $k$ odd lie within observed baby $\mathcal{M}$’s, except that $a_{1}$ and $a_{2n-1}$ lie in the same merged/deformed baby $\mathcal{M}$ near $a=0$ which is not a true baby $\mathcal{M}$. To begin to describe why, we examine the images of the critical value $v_{-}$ under $r_{n,a}(z)$ at the 2n12n-1 $a_{k}$-values listed above.
@@
Thus the n1n-1 points with $k\in\{1,2,3,...,2n-1\},$ for $k$ even we propose are inside of topological disk hyperbolic components and the odd are proposed centers of baby $\mathcal{M}$’s, except in the case of k=1,2n1k=1,2n-1 where the centers are so near 0, the components appear to merge into something more complicated. Hence there appear only to be n2n-2 baby $\mathcal{M}$’s.
@@ 6. Baby Mandelbrots in the $r_{n,a}$ Parameter Space
The location of some baby Mandelbrot sets in the family $R_{n,a,b}$ was established in both [4] and [5]. In the present article we are more interested in the newer phenomena of baby Julia sets in parameter space. So, rather than applying similar techniques to establish the location of baby $\mathcal{M}$’s in this subfamily, we stop at providing some suggestions on how an interested reader could complete this task, by building on the previous articles and the prior results of this article, at least under the assumption $n$ is sufficiently large.
By Lemma 5.1, there are 2n12n-1 potential centers of hyperbolic components, $a_{k}>0,k=1,2,\ldots,2n-1,$ with $v_{-}$ fixed for $k$ odd (n2n-2 values) and $v_{-}$ mapping to $v_{+}$ for $k$ even (n1n-1 values).
@@
To prove this conjecture using Douady and Hubbard’s Theorem 2.9, one must define regions in the $a$-plane $W_{k}$, which are closed topological disks, with with $a_{k}\in W_{k}$, and show that (1): for any $a$ in one of these $W_{k}$’s, the map $r_{n,a}$ restricted to $U^{\prime}_{a,k}$ has a critical point in the $U^{\prime}_{a,k}$, and the map is polynomial-like of degree 2 on the orbits that remain bounded in $U^{\prime}_{a,k}$. So, one needs that $r_{n,a}$ is a degree 2 proper, analytic map of each $U^{\prime}$ to $U^{-}$ where $U^{\prime}$ is contained in $U^{-}$, with $U^{\prime}$ rel. cpt. in $U^{-}$, because by Proposition 2.16, each $U^{\prime}_{a,k}$ maps 2:1 onto to $U^{-}_{a}$ for $k$ odd.
@@
Now, recall $a_{k}=\left(({1-e^{i\frac{k\pi}{n}}})/{4}\right)^{\frac{2n}{n-1}}$. Note we’re excluding $a_{1}$ and $a_{2n-1}$, as those are the parameters that tend to 0 as $n\to\infty$, since $(2n)/(n-1)\to 2$ and $e^{i\pi(2n-1)/n}=e^{-i\pi/n},$ so $(1-e^{\pm i\pi/n})\to 0^{-}$ as $n\to\infty$. Excluding them one has a chance to show the other $W_{k}$’s are bounded away from $a=0$ (and the negative real axis), with a bound depending on $n$. Computer generated images show that there don’t tend to be distinct hyperbolic components for $a_{1}$, and $a_{2n-1}$, rather the components that would correspond to a potential baby $\mathcal{M}$ for $a_{1}$ and $a_{2n-1}$ have “collided”, and they trapped the origin between them, so the two potential baby $\mathcal{M}$’s that cannot be baby $\mathcal{M}$’s, and the component where $D_{-}=D_{+}$ near $a=0$, seem to be deformed into one unusual component. See Figure 4.
One could initially aim to establish this conjecture at least for all $n$ sufficiently large, or even just start with the $W$’s “farthest” from the origin ($a=0$) and negative real axis (the branch cut of $a^{1/2n}$); that is $W_{n}$ if $n$ is odd (which should be roughly centered about the positive real axis) or $W_{n\pm 1},$ if $n$ is even (which are located in symmetrical positions above and below the positive real axis).
One approach to defining the $W_{k}$’s would be to apply the same idea from [4], and define regions $W_{k}$ that should contain baby $\mathcal{M}$s in the $a$-plane by four curves suggested by the definition of $U^{\prime}_{a,k}$. That would yield the following.
@@
However, with the regions defined this way, implicitly as the equations have an $a$ on each side, it complicates some of the remaining criteria. Though, at least for $n$ large, $|a^{1/2n}|$ is close to $1$ (and $1$ was in that spot in the definitions of the $W_{k}$’s in [4]).
For example, note one of the requirements is that each $W_{k}$ be a (closed) topological disk, so with these implicit definitions one would need to check that was true, at least for $n$ sufficiently large.
@@
Also note by Lemma 3.3, from the “outer” argument we see letting $\delta=1$ we get $u=11/16$ and $|v_{-}|>1+\delta=2$ for all $a\in\mathbb{C}\setminus\mathbb{D}(1/8,u=11/16),$ if $n$ sufficiently large, hence inside the curve $\tau$ where $|v_{-}|\leq 2,$ we know our $W_{k}^{\prime}$s are in $\mathbb{D}(1/8,11/16)\subset\mathbb{D}(0,13/16)$ at least for $n$ sufficiently large. So, consider the set $\beta$, where $|v_{-}|=|a|^{1/n}/2$. Now if we take the $n$ sufficiently large that the $W_{k}$’s are in $\mathbb{D}(0,13/16)\subset\mathbb{D}(0,1),$ we have $|a|<1$ on $\beta$, so $|a|^{1/n}/2<1/2$.
@@
Note from, Lemma 3.5 given the upper bound $|v_{-}|=2$ in these $W_{k}$’s we’d have for all $a$ in $W_{k}$, $U^{\prime}_{a,k}\subset\mathcal{E}$.

### v2

@@ Abstract.
11footnotetext: 2020 MSC: Primary: 37F10, 32A20; Secondary: 32A19. Keywords: Complex Dynamical Systems, Mandelbrot, Polynomial-Like, Iteration, Douady, Hubbard
@@ 2. Preliminaries
Notation. Let $\mathcal{N}(S)$ denote a small neighborhood (of unspecified size) of a set $S$.
@@ 2.1. Polynomial-Like Maps
Each polynomial-like map of degree $d$ has exactly d1d-1 critical points. Here we are only concerned with polynomial-like maps of degree two. A polynomial-like map of degree two is a two-to-one map from $U^{\prime}$ to $U$, with a unique critical point in $U^{\prime}$. The “straightening theorem” shows this is an apt name.
@@
Now, since $d=2$ in our case of interest, by the straightening theorem, we can associate each member $f_{\lambda}$ of an analytic family of polynomial-like maps of degree $2$ with a polynomial $P_{\chi(\lambda)}=z^{2}+\chi(\lambda)$ and with a homeomorphism between neighborhoods of their Julia sets
@@
where $\mathcal{N}(K(f_{\lambda}|_{U^{\prime}_{\lambda}}))\subset U^{\prime}_{\lambda}$, and where $\phi_{\lambda}$ is a hybrid equivalence from $f_{\lambda}$ to $P_{\chi(\lambda)}$ (conformal on filled Julia sets, quasi-conformal in neighborhoods), with
@@
We can improve upon this using McMullen and Sullivan’s [24]. In their Chp. 7, they describe the structural stability extended to the Riemann sphere, $\hat{\mathbb{C}}$. Recall a holomorphic family of rational maps $f_{\lambda}$ over a complex manifold $X$ is a holomorphic map $X\times\hat{\mathbb{C}}\to\hat{\mathbb{C}}$, given by $(\lambda,z)\mapsto f_{\lambda}(z).$ In a holomorphic family of rational maps, $f_{\lambda_{0}}$ is topologically stable if for all $\lambda$ in some neighborhood of $\lambda_{0}$, $f_{\lambda}$ and $f_{\lambda_{0}}$ are topologically conjugate (on the sphere). They show this implies there is in fact a quasi-conformal conjugacy, in their Theorem 7.1, and they further show that the set of critical orbit relations is constant.
@@
They use the Harmonic $\lambda$-lemma of Bers-Royden (see [26] and [1]), which shows a holomorphic motion of a set $A$ in a neighborhood in parameter space can be extended uniquely to a holomorphic motion of the entire sphere, over a smaller neighborhood in parameter space (and the motion and its extension agree on the domain overlap). The uniqueness is what allows them to guarantee the compatibility with the dynamics.
In our setting, compatibility with the dynamics implies that the $\tau$’s of Proposition 2.6 and Theorem 2.8 coincide, though Douady and Hubbard’s $\tau_{\lambda}$ was only defined near the Julia sets. This is why we used the same symbol for both of these.
Moreover, McMullen-Sullivan apply their results to hyperbolic quadratic polynomials. Let $H\subset\mathbb{C}$ be the open set of hyperbolic parameters. Partition $H$ by setting $H(\infty)$ to the set of parameters whose critical orbits escape to $\infty$; $H_{0}(p)$ is the set of parameters whose critical point is periodic of order $p$; and $H(p)$ is the set of parameters whose critical orbit is strictly attracted to a cycle of period $p$, so that $H_{0}(p)$ is not in $H(p)$.
Theorem 10.1 ([24]) states first that each connected component of $H(p)$ is isomorphic to a punctured disk and represents a single quasiconformal conjugacy class. The multiplier of the attracting cycle gives a natural isomorphism of $\mathbb{D}^{*}$. Secondly, for each integer $p\geq 1,H_{0}(p)$ is finite, and in fact $H_{0}(p)$ are the punctures in the centers of the components in $H(p)$. In general, one must be careful about centers of hyperbolic components, as $\phi_{\lambda}$ is valid near $J$, and extends via holomorphic motion to all of $\mathbb{C}$ for parameters except hyperbolic component centers. Thirdly, $H(\infty)$ is also isomorphic to a punctured disk and represents a single quasiconformal conjugacy class. Finally, maps in different components of these sets represent different conjugacy classes.
@@
Finally, we give Douady and Hubbard’s criteria for the existence of baby Mandelbrot sets in parameter space, as stated by Devaney based on Douady and Hubbard:
@@ 2.2. Preliminaries Regarding $R_{n,a,b}$
[4] For the family of functions $R_{n,a,b}$, let $\psi=\operatorname{Arg}(a)$ and for each $k=0,\ldots,2n-1,$ define $U_{a,k}^{\prime}$ (which also depends on $n$), as the polar rectangle
@@
So $U^{\prime}_{a,k}$ is the polar rectangle with arguments centered at $(\psi+2k\pi)/2n$ and then plus or minus $\pi/2n$. Note $U^{\prime}_{a,k}$ does not depend on $b$, but its image does.
@@
Note before the shift and rotation, the ellipse $\mathcal{E}$ is centered at 0, with its semi-major axis lying along the $x$-axis and of length $\left(2^{n}+{|a|}/{2}\right)$, and its semi-minor axis lying along the $y$-axis and of length $\left(2^{n}-{|a|}/{2}\right)$.
@@
[4] For each $k$ in $\{0,1,\ldots,2n-1\}$, if $k$ is even, $U_{a}^{+}=R_{n,a,b}(U^{\prime}_{a,k})$, and this set is one half of the ellipse $\mathcal{E}$ including the minor axis and the critical value $v_{+}$. Moreover, $R_{n,a,b}$ maps $U^{\prime}_{a,k}$ $2:1$ onto $U_{a}^{+}$ for each even $k$.
@@ 3. Dynamical Plane Results for $r_{n,a}$
Excluding a “degenerate” case. For $n\geq 3$, if $a$ is small it can happen that $D_{-}=D_{+}$, that is, $v_{-}$ and $v_{+}$ are in the same (bounded) Fatou component. In that case, there is only one Fatou component, which the reader may imagine as a topological disk with an infinite tree of disks removed (the Fatou component containing $z=0$ and all of its preimages). Xiao, Qiu, and Yin ([27]) call such a $J$ a Cantor set of bubbles. In our subfamily with one critical point fixed, the hyperbolic component containing $a=0$ is a quite small teardrop-like shape aligned along the real axis, roughly in the bounds from $\Re(a)=-0.004$ to $\Re(a)=0.0025$ and $\Im(a)=-0.0025$ to $\Im(a)=0.0025.$ In the images in Figure 4, this component is not really visible as it lies just to the right of the “deformed” baby $\mathcal{M}$-like set near the cusp of the cardioid shape of the bifurcation locus. But in the general, parameter space maps with $D_{-}=D_{+}$ are quite easy to find. For example, on the left of Figure 6 is the parameter slice which is the $b$-plane, holding $a$ small and constant, in this case for $n=4,a=0.0001+0.000i.$ The image appears like the parameter space for $z^{4}+b$, except with “bites” taken out of it due to the singular perturbation of 0 mapping to $\infty$. Any point in the main central annulus-like black component has both critical values in this common Fatou component. One such example is in this figure on the right, for $n=3,a=0.001+0.001i,c=0.25+0.4i$, although the situation for $n\geq 3$ is similar. The critical values are marked, and the shading inside of the filled Julia set illustrates the rate of attraction to the attracting fixed point.
@@ 4. Locating the bifurcation locus of $r_{n,a}$
We next show the bifurcation locus lies within an annulus. In the limit of the above cardioid as $n\to\infty$, the cardioid has cusp at $a(\pi)=0$, and max on the $x$-axis at $a(0)=1/4$, so is contained in an annulus centered at $1/8$, with a small inner radius <1/8<1/8, and outer radius about $1/7$. We show below is that the boundedness locus is contained in the annulus $\mathbb{A}(l,u)+1/8$, with center $1/8$, outer radius any $u>3/8$ and inner radius any $l\leq 1/32,$ for $n$ sufficiently large.
@@
One way to improve on Lemma 4.2 would be to establish the following:
@@
The outline of a proof could be somewhat similar to the proof of Theorem 2 given in [4] (or to the proof of a simpler case, for a lemma in [6]): show that inside the annulus guaranteed by Lemma 4.2, but outside of $\mathcal{N}_{\varepsilon}(\mathcal{S}_{n}),$ $|v_{-}|$ is bounded away from $1$, then apply Corollary 2.10. The fact that the spine depends on $n$ is an added complication that does not appear in [4], but computer images suggest $M_{n}(r_{n,a})\to\mathcal{S}_{\infty}$ as $n\to\infty$.
@@ 5. Centers of Principal Components in the $r_{n,a}$ Parameter Space
This gives us one trivial solution: If $a=0$, then $b-2\sqrt{a}=b+2\sqrt{a}=a^{1/2n}=0$. For the rest, we need $v_{-}$ to be equal to a non-canonical choice of root.
@@
Around these 2n12n-1 points, we see two patterns emerge. Parameters defined by the above with $k$ even lie inside the disks we see in the bifurcation locus. The $a_{k}$ for $k$ odd lie within observed baby $\mathcal{M}$’s, except that $a_{1}$ and $a_{2n-1}$ lie in the same merged/deformed baby $\mathcal{M}$ near $a=0$, which is not a true baby $\mathcal{M}$. To begin to describe why, we examine the images of the critical value $v_{-}$ under $r_{n,a}(z)$ at the 2n12n-1 $a_{k}$-values listed above.
@@
Thus the n1n-1 points with $k\in\{1,2,3,...,2n-1\},$ for $k$ even we propose are inside of topological disk hyperbolic components, and the points with $k$ odd are proposed centers of baby $\mathcal{M}$’s, except in the case of k=1,2n1k=1,2n-1, where the centers are so near 0, the components appear to merge into something more complicated. Hence there appear only to be n2n-2 baby $\mathcal{M}$’s.
@@ 6. Baby Mandelbrots in the $r_{n,a}$ Parameter Space
The locations of some baby Mandelbrot sets in the family $R_{n,a,b}$ were established in both [4] and [5]. In the present article we are more interested in the newer phenomena of baby Julia sets in parameter space. So, rather than applying similar techniques to establish the location of baby $\mathcal{M}$’s in this subfamily, we stop at providing some suggestions on how an interested reader could complete this task, by building on the previous articles and the prior results of this article, at least under the assumption that $n$ is sufficiently large.
By Lemma 5.1, there are 2n12n-1 potential centers of hyperbolic components, n2n-2 values of the form $a_{k}>0,k=1,2,\ldots,2n-1,$ with $v_{-}$ fixed for $k$ odd, and n1n-1 values with $v_{-}$ mapping to $v_{+}$ for $k$ even.
@@
To prove this conjecture using Douady and Hubbard’s Theorem 2.9, one must define regions $W_{k}$ in the $a$-plane, which are closed topological disks with $a_{k}\in W_{k}$, and show that (1): for any $a$ in one of these $W_{k}$’s, the map $r_{n,a}$ restricted to $U^{\prime}_{a,k}$ has a critical point in the $U^{\prime}_{a,k}$, and the map is polynomial-like of degree 2 on the orbits that remain bounded in $U^{\prime}_{a,k}$. So, one needs that $r_{n,a}$ is a degree 2 proper, analytic map of each $U^{\prime}$ to $U^{-}$ where $U^{\prime}$ is contained in $U^{-}$, with $U^{\prime}$ relatively compact in $U^{-}$, because by Proposition 2.16, each $U^{\prime}_{a,k}$ maps 2:1 onto to $U^{-}_{a}$ for $k$ odd.
@@
Now, recall $a_{k}=\left(({1-e^{i\frac{k\pi}{n}}})/{4}\right)^{\frac{2n}{n-1}}$. Note we’re excluding $a_{1}$ and $a_{2n-1}$, as those are the parameters that tend to 0 as $n\to\infty$, since $(2n)/(n-1)\to 2$ and $e^{i\pi(2n-1)/n}=e^{-i\pi/n},$ so $(1-e^{\pm i\pi/n})\to 0^{-}$ as $n\to\infty$. Excluding them one has a chance to show the other $W_{k}$’s are bounded away from $a=0$ (and the negative real axis), with a bound depending on $n$. Computer generated images show that there don’t tend to be distinct hyperbolic components for $a_{1}$, and $a_{2n-1}$, but rather that the components that would correspond to potential baby $\mathcal{M}$s for $a_{1}$ and $a_{2n-1}$ have “collided”, with the origin trapped between them. If this is the case, then the two potential baby $\mathcal{M}$s cannot be baby $\mathcal{M}$s, and the component where $D_{-}=D_{+}$ near $a=0$ seems to be deformed into one unusual component. See Figure 4.
One could initially aim to establish this conjecture at least for all $n$ sufficiently large, or even just start with the $W$s “farthest” from the origin (where $a=0$) and negative real axis, which is the branch cut of $a^{1/2n}$; that is, one could start with $W_{n}$ if $n$ is odd, which should be roughly centered about the positive real axis, or with $W_{n\pm 1},$ if $n$ is even, which are located in symmetrical positions above and below the positive real axis.
One approach to defining the $W_{k}$’s would be to apply the same idea from [4] and define regions $W_{k}$ that should contain baby $\mathcal{M}$s in the $a$-plane by using four curves suggested by the definition of $U^{\prime}_{a,k}$. That would yield the following.
@@
On the other hand, the region definitions given above are implicit, as the equations have an $a$ on each side, which complicates some of the remaining criteria. However, at least for $n$ large, $|a^{1/2n}|$ is close to $1$, and $1$ was in that spot in the definitions of the $W_{k}$’s in [4], so this may not be too difficult of a complication.
A place where this implicit definition may complicate the process is in the requirement that each $W_{k}$ be a (closed) topological disk, as one would need to check that was true, at least for $n$ sufficiently large.
@@
Also, note by Lemma 4.2, from the “outer” argument, we see that in letting $\delta=1$, we get $u=11/16$ and so for all $n$ sufficiently large and for all $a\in\mathbb{C}\setminus\mathbb{D}(1/8,u=11/16),$ we get that $|v_{-}|>1+\delta=2$. Thus, inside the curve $\tau$ where $|v_{-}|\leq 2,$ we know our $W_{k}^{\prime}$s are in $\mathbb{D}(1/8,11/16)\subset\mathbb{D}(0,13/16)$ at least for $n$ sufficiently large. So, consider the set $\beta$, where $|v_{-}|=|a|^{1/n}/2$. Now if we take the $n$ sufficiently large that the $W_{k}$’s are in $\mathbb{D}(0,13/16)\subset\mathbb{D}(0,1),$ we have $|a|<1$ in $\beta$, so $|a|^{1/n}/2<1/2$.
@@
Note from, Lemma 3.5 given the upper bound $|v_{-}|=2$ in these $W_{k}$’s we’d have $U^{\prime}_{a,k}\subset\mathcal{E}$ for all $a$ in $W_{k}$.